### v1

@@ 1 Introduction
The Giant Radio Array for Neutrino Detection (GRAND) [1, 2] is a planned detector of ultra-high-energy (UHE) cosmic particles, with energies exceeding $10^{17}$ eV. Its primary goal is to find the long-sought origin of UHE cosmic rays (UHECRs), purportedly the most energetic cosmic accelerators [3]. GRAND will achieve this directly by detecting large numbers of UHECRs and, indirectly, by being sensitive to the potentially tiny fluxes of associated ultra-high-energy (UHE) gamma rays and, especially, neutrinos, first predicted in 1969 [4, 5].
GRAND uses sparse, vast ground arrays of antennas to detect the radio emission from extensive air showers triggered by UHE particles [6, 7] that interact in the atmosphere or just below the surface, including in surrounding mountains. Because radio waves have a long attenuation length in air, the arrays monitor large volumes of air and ground, conferring GRAND the large exposure needed to detect UHE particles even if their fluxes are low, and the sub-degree angular resolution needed to pinpoint sources of UHE neutrinos [5, 8, 9]. When looking for UHE neutrinos, GRAND targets air showers initiated by Earth-skimming tau neutrinos, which, on average, survive traveling underground more often than electron and muon neutrinos [10].
GRAND has a staged construction plan, each stage involving progressively larger arrays [1, 2]. An array is made up of detection units (DUs), each consisting of a passive antenna that detects the radio emission from air showers and electronics to amplify, filter, and digitize it. In its final stage, GRAND is envisioned as a collection of geographically separate arrays containing thousands of DUs each [1, 2].
Presently, three prototype GRAND arrays are in operation  [11, 2, 12]: GRANDProto300 in Dunhuang, China, with its first 13 DUs out of 300 running, to be expanded into a 200-km2 array in the near future; GRAND@Auger at the site of the Pierre Auger Observatory in Malargüe, Argentina, with 10 DUs; and GRAND@Nançay, at the Nançay Radio Observatory, in France, with 4 DUs. Their goal is twofold: to test the performance and deployment of the GRAND DUs under field conditions and to validate the autonomous detection of extensive air showers—i.e., large particle showers that develop in the atmosphere—reaching the array from near-horizontal directions, as expected from UHE neutrinos.
@@
In addition, GRAND is expected to have a high throughput of data. We estimate that GRANDProto300 alone will generate 1.1 petabytes (PB) of data during its construction, of which 400 terabytes (TB) will be simulations, 450 TB will be data collected by the array, and 250 TB will be produced by post-processing analyses. Once GRANDProto300 is completed, we estimate the data volume to grow to 2.5 PB per year. Extracting physical insight from such a large volume of data requires software that is able to analyze experimental data and simulate the generation, propagation, and detection of the radio signals from extensive air showers.
@@
Figure 1: Overview of the major components of GRANDlib. The simulation pipeline of GRANDlib mimics closely that of the real hardware and software capabilities of a radio-detection array like GRAND [1]. An ultra-high-energy (UHE) particle triggers an extensive air shower whose radio emission [6, 7] is picked up by an array of detection units at ground level. In a detection unit, an antenna receives the incident radio signal and converts it into an open-circuit voltage, to which GRANDlib adds noise. The resulting voltage signal is then processed through the radio-frequency chain of the unit for amplification—yielding the final output voltage—digitization, and storage. Beside the simulation pipeline, GRANDlib has tools to handle coordinate systems, geomagnetic field, the topography of Earth, and data management. GRANDlib can be installed from Ref. [19], which also contains usage examples.
@@ 2 Overview of GRANDlib
GRANDlib is designed to generate simulations of the radio-detection of extensive air showers in GRAND, and to assess the effect of changes to the detector design on simulated observations, while meeting the requirements for simulation production, signal processing, data storage, and data analysis.
GRANDlib uses descriptions of the antenna response and the radio-frequency (RF) chain—the series of electronic components that process and digitize the signal received by an antenna while minimizing background noise—that are specific to GRAND DUs, but has built-in flexibility to accommodate other, similar experiments. Some GRANDlib modules are particularly useful to other experiments, especially those that deal with coordinate systems, topography, and the geomagnetic field.
@@
Shower generation (Section 6) is outsourced to simulation tools external to GRANDlib. DANTON [13] handle the simulation of taus made in $\nu_{\tau}$ interactions underground, and ZHAireS [15] or CoREAS [16] handles the simulation of extensive air showers and the radio signals from them.
@@
GRANDlib shares some common functionality with simulation software custom-designed for other experiments that detect UHE particles via their radio emission in air [20]—like the Pierre Auger Observatory [21]—or in ice [22, 23]—like RNO-G [24] and the planned IceCube-Gen2 [25].
Table 1 lists the units of the quantities used in GRANDlib, including in the code examples shown in Listings 14.
Quantity Unit
Energy Gigaelectronvolt (GeV)
Time Nanosecond (ns)
Length Meter (m)
Voltage Microvolt ($\mu$V)
Electric field Microvolt per meter ($\mu$V m-1)
Magnetic field Tesla (T)
Angle Degree
Grammage g cm-2
Density g cm-3
Frequency Hertz (Hz)
Resistance Ohm ($\Omega$)
Table 1: Units of physical quantities used in GRANDlib.
@@ 3 Coordinate systems
The GRANDlib coordinates module handles the different coordinate systems used by GRAND: Earth-centered Earth-fixed (ECEF), geodetic, Local Tangential Plane (LTP), and GRANDCS, the primary coordinate system used in GRAND. A contains a detailed description of the systems and their transformation from one into another. All coordinate systems account for the curvature of the Earth. [In the simulations run to produce the results shown in this paper, the antenna positions were placed on a flat Earth—other than accounting for differences in ground elevation—which, given the relatively small size of the array, amounts to only a small error.]
@@
3# Define coordinate point/s. Coordinates can be a number or an array. Unit of x,y,z, and height is in meters. Unit of longitude and latitude is in degrees.
4ecef = ECEF(x=6378137, y=0, z=0) # at the equator
5# At the GP300 site at the sea level
6geod = Geodetic(latitude=40.98, longitude=93.95, # deg
7 height=0) # m
8# At the GP300 site on the ground
9GP_loc = Geodetic(latitude=40.98, longitude=93.95, # deg
10 height=1267.0) # m
11ltp = LTP(x=10, y=10, z=10, # x, y, z are in meters
12 location=geod, orientation=’NWU’)
13ltp_GP = LTP(x=10, y=10, z=10, # x, y, z are in meters
14 location=GP_loc, orientation=’NWU’)
15grandcs = GRANDCS(x=10, y=10, z=10, location=GP_loc)
16
17# These coordinate systems (CS) can be transformed among themselves by two easy-to-use methods:
18# Method 1. Instantiating a required CS with a given CS.
19# Method 2. Calling a method from the given CS object.
20# Only a code snippet for method 1 is shown here.
21
22# Convert Geodetic, LTP, and GRANDCS to ECEF resp.
23ecef1,ecef2,ecef3 = ECEF(geod),ECEF(ltp),ECEF(grandcs)
24print(’ecef1’, ecef1, ’m’)
25print(’ecef2’, ecef2, ’m’)
26print(’ecef3’, ecef3, ’m’)
27
28# Convert ECEF, LTP, and GRANDCS to Geodetic resp.
29geod1 = Geodetic(ecef, reference="ELLIPSOID")
30geod2, geod3 = Geodetic(ltp), Geodetic(grandcs)
31print(’geod1’, geod1, ’deg,deg,m’)
32print(’geod2’, geod2, ’deg,deg,m’)
33print(’geod3’, geod3, ’deg,deg,m’)
34
35# Convert to LTP.
36ltp1 = LTP(ecef, location=GP_loc, orientation=’NWU’)
37ltp2 = LTP(geod, location=GP_loc, orientation=’NWU’)
38ltp3 = LTP(grandcs, location=GP_loc, orientation=’NWU’)
39ltp4 = LTP(ltp, location=GP_loc, orientation=’NWU’)
40print(’ltp1’, ltp1, ’m’)
41print(’ltp2’, ltp2, ’m’)
42print(’ltp3’, ltp3, ’m’)
43print(’ltp4’, ltp4, ’m’)
44
45# Convert ECEF, Geodetic, and LTP to GRANDCS resp.
46gdcs1, gdcs2, gdcs3 = GRANDCS(ecef), GRANDCS(geod), GRANDCS(ltp)
47print(’grandcs1’, gdcs1, ’m’)
48print(’grandcs2’, gdcs2, ’m’)
49print(’grandcs3’, gdcs3, ’m’)
50
51# This script returns:
52# ecef1 [[-332170.75], [4810587.48], [4160741.60]] m
53# ecef2 [[-332160.84], [4810589.16], [4160755.71]] m
54# ecef3 [[-332226.77], [4811543.36], [4161586.64]] m
55
56# geod1 [[0.], [0.], [0.]] # deg, deg, m
57# geod2 [[40.98009],[93.94988],[9.99957]] deg,deg,m
58# geod3 [[40.98009],[93.94988],[1276.99957]] deg,deg,m
59
60# ltp1 [[309302.28], [6362986.02], [-6701900.83]] m
61# ltp2 [[0.], [0.], [-1267.00]] m
62# ltp3 [[10.05], [9.95], [ 10.0]] m
63# ltp4 [[10.0], [10.0], [-1257.0]] m
64
65# grandcs1 [[118397.35], [6374520.29], [-6570655.99]] m
66# grandcs2 [[234807.77], [-137733.72], [-8739.44]] m
67# grandcs3 [[234817.85], [-137723.66], [-8729.58]] m
Listing 1: GRANDlib code snippet to convert among different coordinate systems. GRANDCS is the primary coordinate system used by GRAND. See Section 3 for details. More examples are available in the GRANDlib repository [19], under examples/geo/coordinates.ipynb.
@@ 4 Topography
The topography module handles the topography of the Earth, which determines, for instance, where an UHE $\nu_{\tau}$ could interact underground to make a tau that initiates a shower. It is based on TURTLE (Topographic Utilities for tRansporting parTicules over Long rangEs) [14, 26], which provides utilities for simulating the long-range transport of particles through a pre-defined topography. TURTLE uses ray-tracing to follow particle trajectories; the calculation is sped up by using only the topography neighboring the particle along its trajectory.
The topography module computes the ground elevation at locations on the surface of the Earth from 56S to 60N latitude. It uses NASA Shuttle Radar Topography Mission (SRTM) [27, 28] data with a resolution of 1 arcsecond. Elevation can be computed in reference to the sea level, the ellipsoidal Earth model, or a user-defined LTP frame (Section 3). Limited topographical data is included in the GRANDlib installation; users can download more from NASA Earthdata [29].
Figure 2 shows, as an example, the topographic map of the area surrounding the GRANDProto300 site. Mountainous regions like the one in Fig. 2 are ideal locations for GRAND, since the mountains act as possible targets for neutrino interactions [10, 1].
@@
Figure 2: Topography of the area surrounding the GRANDProto300 prototype array in Dunhuang, China. The array is centered at 40.98 latitude and 93.95 longitude. Overlaid is one of the proposed layouts of GRANDProto300; see Fig. 9. Locations with surrounding mountains present an ideal topography for GRAND sites. The arrow illustrates the trajectory of a particle starting in air and hitting a nearby mountain; see Fig. 3. The topography data used here is from the NASA SRTM mission. See Section 4 for details.
@@
1# Code for Figures 2 and 3.
2from grand import topography
3import numpy as np
4
5# Build a meshgrid for ground topography shown in Fig.2.
6xmin, xmax, ymin, ymax = -5e4, +5e4, -5e4, +5e4
7x = np.linspace(xmin, xmax, 401)
8y = np.linspace(ymin, ymax, 401)
9X, Y = np.meshgrid(x, y)
10# Compute elevation around the GP300 site.
11geod = Geodetic(latitude=40.98, longitude=93.95,
12 height=0) # at the sea level
13grnd_grid = GRANDCS(x=X.flatten(), y=Y.flatten(),
14 z=np.zeros(X.size), location=geod)
15
16# Calculate elevation wrt geoid. This is shown in Fig 2.
17zg = topography.elevation(grnd_grid, "GEOID")
18zg = zg.reshape(X.shape) # Format as matrix
19
20# Now go to Geodetic coordinates
21grnd_geod = Geodetic(grnd_grid)
22
23# Now reformat output as matrix
24lat = grnd_geod.latitude.reshape(X.shape)
25lat = lat[0] # Grab first line
26lon = grnd_geod.longitude.reshape(X.shape)
27lon = lon[:, 0] # Grab first col
28
29# direction vector of a moving particle in GRANDCS ref
30direction = GRANDCS(x=10, y=-10, z=-0.036,
31 location=geod)
32# Initial point along traj (in GRANDCS ref)
33x0 = GRANDCS(x=-4e4, y=4e4, z=1850, location=geod)
34# GRANDCS --> ECEF. input direction must be in ECEF.
35dirECEF = np.matmul(x0.basis.T, direction)
36dist = topography.distance(x0, dirECEF)
37print("Distance to ground:", dist, "m")
38
39# Build trajectory of a moving particle.
40u = np.linspace(0, 1.1*dist, 401) # distance from x0 (m)
41# normalize direction vector
42dirn = direction / np.linalg.norm(direction)
43traj = dirn * u + x0
44
45# Get elevation below traj. This is shown in Figure 3.
46traj_grand = GRANDCS(x=traj[0], y=traj[1], z=traj[2], location=geod) # Compute traj coordinates
47# z-coordinate of ground in GRANDCS ref
48ztG = topography.elevation(traj_grand,reference="LOCAL")
49# index of trajectory closest to intersection to ground
50icrash = np.argmin(abs(u - dist))
51
52# This scripts gives:
53# "zg" is the elevation shown in Figure 2.
54# "ztG" is the elevation shown in Figure 3. Crash point is given by "u[icrash], traj[2, icrash]"
55# Distance to ground: 37374.39 m
@@ 5 Geomagnetic field
1# Magnetic field information at the GP300 site used to generate the simulations for this paper.
@@
4# Compute geomagnetic field at GP300 site.
5GP_loc = Geodetic(latitude=40.98,longitude=93.95, # deg
6 height=1267.0) # m
7
8# Geomagnetic field at the GP300 location.
9# Accepted formats for ‘obstime’ are ‘2020-01-01’ and datetime.date(2020, 1, 1).
10geoB = Geomagnet(location=GP_loc, obstime="2022-08-28")
11
12print("Mag. Field (B):", geoB.field, "Tesla")
13print("Observation time:", geoB.obstime)
14print("Location:", geoB.location)
15print("Declination [deg]:", round(geoB.declination, 3))
16print("Inclination [deg]:", round(geoB.inclination, 3))
17
18# This script returns:
19# Mag. Field (B):
20# [[4.4571e-08], [2.6575e-05], [-4.9974e-05]] Tesla
21# Observation time: 2022-08-28
22# Location: [[40.98], [93.95], [1267.0]]
23# Declination [deg]: 0.096
24# Inclination [deg]: 61.997

### v2

@@ Abstract
keywords:
GRAND , astroparticle , ultra-high energy physics , neutrino , cosmic ray , experiment
@@ 1 Introduction
The Giant Radio Array for Neutrino Detection (GRAND) [1, 2, 3] is a planned detector of ultra-high-energy (UHE) cosmic particles, with energies exceeding $10^{17}$ eV. Its primary goal is to find the long-sought origin of UHE cosmic rays (UHECRs), purportedly the most energetic cosmic accelerators [4]. GRAND will achieve this directly by detecting large numbers of UHECRs and, indirectly, by being sensitive to the potentially tiny fluxes of associated ultra-high-energy (UHE) gamma rays and, especially, neutrinos, first predicted in 1969 [5, 6].
GRAND uses sparse, vast ground arrays of antennas to detect the radio emission from extensive air showers triggered by UHE particles [7, 8] that interact in the atmosphere or just below the surface, including in surrounding mountains. Because radio waves have a long attenuation length in air, the arrays monitor large volumes of air and ground, conferring GRAND the large exposure needed to detect UHE particles even if their fluxes are low, and the sub-degree angular resolution needed to pinpoint sources of UHE neutrinos [6, 9, 10]. When looking for UHE neutrinos, GRAND targets air showers initiated by Earth-skimming tau neutrinos ($\nu_{\tau}$), which, on average, survive traveling underground more often than electron and muon neutrinos [11].
GRAND has a staged construction plan, each stage involving progressively larger arrays [1, 2, 3]. An array is made up of detection units (DUs), each consisting of a passive antenna that detects the radio emission from air showers and electronics to amplify, filter, and digitize it. In its final stage, GRAND is envisioned as a collection of geographically separate arrays containing thousands of DUs each [1, 2].
Presently, three prototype GRAND arrays are in operation  [12, 2, 13, 3, 14]: GRANDProto300 in Dunhuang, China, with its first 48 DUs out of 300 deployed, to be expanded into a 200-km2 array in the near future; GRAND@Auger at the site of the Pierre Auger Observatory in Malargüe, Argentina, with 10 DUs; and GRAND@Nançay, at the Nançay Radio Observatory, in France, with 4 DUs. Their goal is twofold: to test the performance and deployment of the GRAND DUs under field conditions and to validate the autonomous detection of extensive air showers—i.e., large particle showers that develop in the atmosphere—reaching the array from near-horizontal directions, as expected from UHE neutrinos.
@@
In addition, GRAND is expected to have a high throughput of data. We estimate that GRANDProto300 alone will generate 1.1 petabytes (PB) of data during its construction, of which 400 terabytes (TB) will be simulations, 450 TB will be collected by the array, and 250 TB will be produced by post-processing analyses. Once GRANDProto300 is completed, we estimate the data volume to grow to 2.5 PB per year. [Out of this, 2 PB comes from an estimated event rate of 100 Hz, 20 antennas triggered per event (in an event, the radio signal fulfills pre-defined trigger conditions), and 33 kB of data created per antenna hit. There is an additional 0.5 PB coming from GRAND simulations and analyses.] Extracting physical insight from such a large volume of data requires software that is able to analyze experimental data and simulate the generation, propagation, and detection of the radio signals from extensive air showers.
@@
Figure 1: Overview of the major components of GRANDlib. The simulation pipeline of GRANDlib mimics closely that of the real hardware and software capabilities of a radio-detection array like GRAND [1]. An ultra-high-energy (UHE) particle triggers an extensive air shower whose radio emission [7, 8] is picked up by an array of detection units at ground level. In a detection unit, an antenna receives the incident radio signal and converts it into an open-circuit voltage, to which GRANDlib adds noise. The resulting voltage signal is then processed through the radio-frequency chain of the unit for amplification—yielding the final output voltage—digitization, and storage. Beside the simulation pipeline, GRANDlib has tools to handle coordinate systems, geomagnetic field, the topography of Earth, and data management. GRANDlib can be installed from Ref. [21], which also contains usage examples. (The GRAND hardware includes an impedance-matching network between the antenna and the low-noise amplifier, not pictured in the radio-frequency chain presented here, but present in an imminent GRANDlib release.)
@@ 2 Overview of GRANDlib
GRANDlib is designed to generate simulations of the radio-detection of extensive air showers in GRAND, and to assess the effect of changes to the detector design on simulated observations, while meeting the requirements for simulation production, signal processing, data storage, and data analysis. To achieve this efficiently, GRANDlib interfaces several existing external packages, as described below.
GRANDlib uses descriptions of the antenna response and the radio-frequency (RF) chain—the series of electronic components that process and digitize the signal received by an antenna while minimizing background noise—that are specific to GRAND DUs, but has built-in flexibility to accommodate other, similar experiments.
@@
Shower generation (Section 6) is outsourced to simulation tools external to GRANDlib. DANTON [15] handles the simulation of taus made in $\nu_{\tau}$ interactions underground, and ZHAireS [17] or CoREAS [18] handles the simulation of extensive air showers and the radio signals from them.
@@
GRANDlib is built on the basis of software developed for TREND [22]. It shares some common functionality with simulation software custom-designed for other experiments that detect UHE particles via their radio emission in air [23]—like the Pierre Auger Observatory [24]—or in ice [25, 26]—like RNO-G [27] and the planned IceCube-Gen2 [28]. It benefits from experience gained with well-established software packages developed by other experiments, like the IceCube Neutrino Observatory [29] and the Pierre Auger Observatory [30]. Beyond its primary applications, GRANDlib can serve other, similar experiments as a user-friendly tool to convert between shower and array coordinate systems, perform topography-related computations, and query the value of the geomagnetic field at different locations; in addition, it offers a data format optimized for efficient storage and processing.
@@ 3 Coordinate systems
The GRANDlib coordinates module handles the different coordinate systems used by GRAND: Earth-centered Earth-fixed (ECEF), geodetic, Local Tangential Plane (LTP), and GRANDCS, the primary coordinate system used in GRAND. A contains a detailed description of the systems and their transformation from one into another. All coordinate systems account for the curvature of the Earth.
@@
3# At the GP300 site
4geod = Geodetic(latitude=40.98, longitude=93.95, # deg
5 height=1267) # m
6gdcs = GRANDCS(x=10, y=10, z=10, location=geod) # m
7
8# Convert GRANDCS to ECEF, Geodetic, and LTP
9gd2ecef, gd2geod = ECEF(gdcs),Geodetic(gdcs)
10gd2ltp = LTP(gdcs, location=gdcs, orientation=’NWU’)
11print(’gd2ecef’, gd2ecef, ’m’)
12print(’gd2geod’, gd2geod) # in deg,deg,m
13print(’gd2ltp’, gd2ltp, ’m’)
14
15# This script returns:
16# gd2ecef [[-332226.78], [4811543.36], [4161586.64]] m
17# gd2geod [[40.98009048], [93.9498818], [1276.9996]]
18# gd2ltp [[0], [0], [0]] m
Listing 1: GRANDlib code snippet to convert GRANDCS to different coordinate systems. GRANDCS is the primary coordinate system used by GRAND. See Section 3 for details. More examples are available in the GRANDlib repository [21], under examples/geo/coordinates.ipynb.
@@ 4 Topography
The topography module handles the topography of the Earth, which determines, for instance, where an UHE $\nu_{\tau}$ could interact underground to make a tau that initiates a shower. It is based on TURTLE (Topographic Utilities for tRansporting parTicules over Long rangEs) [16, 31], which provides utilities for simulating the long-range transport of particles through a pre-defined topography. TURTLE uses ray tracing to follow particle trajectories; the calculation is sped up by using only the topography neighboring the particle along its trajectory.
The topography module computes the ground elevation at locations on the surface of the Earth from 56S to 60N latitude. It uses NASA Shuttle Radar Topography Mission (SRTM) [32, 33] data with a resolution of 1 arcsecond; along the equator, this corresponds to a resolution of about 30 m. Elevation can be computed in reference to the sea level, the ellipsoidal Earth model, or a user-defined LTP frame (Section 3). Limited topographical data is included in the GRANDlib installation; users can download more from NASA Earthdata [34].
Figure 2 shows, as an example, the topographic map of the area around the GRANDProto300 site. Mountainous regions like in Fig. 2 are ideal locations for GRAND, since the mountains act as possible targets for neutrino interactions [11, 1].
@@
1from grand import topography
2import numpy as np
3
4# GRANDCS at the GP300 site defined asl
5geod_asl = Geodetic(latitude=40.98, longitude=93.95,#deg
6 height=0) # m
7
8# Build a meshgrid for ground topography shown in Fig.2.
9bound = np.linspace(-5e4, +5e4, 401)
10X, Y = np.meshgrid(bound, bound)
11grnd_grid = GRANDCS(x=X.flatten(), y=Y.flatten(),
12 z=np.zeros(X.size), location=geod_asl)
13
14# Calculate elevation wrt geoid. This is shown in Fig.2.
15zg = topography.elevation(grnd_grid, "GEOID")
16
17# Initial point along traj (in GRANDCS ref)
18x0 = GRANDCS(x=-4e4, y=4e4, z=1850, location=geod_asl)
19# direction vector of a moving particle in GRANDCS ref
20direction = GRANDCS(x=10, y=-10, z=-0.036,
21 location=geod_asl)
22# Direction vector must be in ECEF.
23dirECEF = np.matmul(x0.basis.T, direction)
24dist = topography.distance(x0, dirECEF)
25print("Distance to ground:", dist, "m")
26
27# Build trajectory of a moving particle.
28u = np.linspace(0, 1.1*dist, 401) # distance from x0 (m)
29# normalize direction vector
30dirn = direction / np.linalg.norm(direction)
31traj = dirn * u + x0
32# Get elevation below traj. This is shown in Fig.3.
33traj_grand = GRANDCS(x=traj[0], y=traj[1], z=traj[2], location=geod_asl) # Compute traj coordinates
34# z-coordinate of ground in GRANDCS ref
35ztG = topography.elevation(traj_grand,reference="LOCAL")
36
37# This scripts gives:
38# "zg" is the elevation shown in Fig.2.
39# "ztG" is the elevation shown in Fig.3.
40# Distance to ground: 37374.39 m
Figure 2: Topography of the area surrounding the GRANDProto300 prototype array in Dunhuang, China. The array is centered at 40.98 latitude and 93.95 longitude. Overlaid is one of the proposed layouts of GRANDProto300; see Fig. 9. Locations with surrounding mountains present an ideal topography for GRAND sites. The arrow illustrates the trajectory of a particle starting in air and hitting a nearby mountain; see Fig. 3. The topography data used here is from the NASA SRTM mission. See Section 4 for details. (For interpretation of the colors in the figure(s), the reader is referred to the web version of this article.)
@@ 5 Geomagnetic field
1# Geomagnetic field information at the GP300 site used to generate the simulations for this paper.
@@
4# Accepted formats for ‘obstime’ are ‘2020-01-01’ and datetime.date(2020, 1, 1).
5geoB = Geomagnet(location=geod, obstime="2022-08-28")
6
7print("Mag. Field (B):", geoB.field, "Tesla")
8print("Observation time:", geoB.obstime)
9print("Location:", geoB.location)
10print("Declination [deg]:", round(geoB.declination, 3))
11print("Inclination [deg]:", round(geoB.inclination, 3))
12
13# This script returns:
14# Mag. Field (B):
15# [[4.4571e-08], [2.6575e-05], [-4.9974e-05]] Tesla
16# Observation time: 2022-08-28
17# Location: [[40.98], [93.95], [1267.0]]
18# Declination [deg]: 0.096
19# Inclination [deg]: 61.997